OpenStudy (anonymous):

Suppose that a department contains 13 men and 18 women. How many ways are there to form a committee with 6 members if it must have strictly more women than men?

OpenStudy (jack1):

yo, minato, howzit dude?!

OpenStudy (anonymous):

That's Hokage-Sama to you (jk xp)

OpenStudy (jack1):

committee must have more women than men so x + y = 6 y > x so y = 4, 5 or 6 sidebar: ...baka, think I'll call u jii-jii! ;D

OpenStudy (jack1):

anyways, back to the Q ;)

OpenStudy (anonymous):

I appreciate your assistance... *bows deeply*

OpenStudy (jack1):

committee must have more women than men so x + y = 6 y > x so y = 4, 5 or 6 im not great at combinatorix, but i'll give it a go: so for the men: options are 2,1 and 0 binomial expansion of \[\frac nm = \frac {13}{2} ~and~ \frac {13}{1} and = 78 ~and~ 13 \]

OpenStudy (jack1):

for the women: options are 4,5 and 6 \[\frac nm = \frac {18}{4} ~and~ \frac {18}{5}~ and~\frac {18}{6} = 3060 ~and~ 8568 ~and~ 18564 \] so 78*3060 + 13* 8568 + 18564 = final answer... i think... but i'm not certain whether that's too specific: ie cares if they're in a certain order?

OpenStudy (jack1):

hmm... @hartnn ...gotta sec to check over this please?

OpenStudy (anonymous):

The question doesn't detail it more than that OnO

OpenStudy (anonymous):

You were right.

OpenStudy (jack1):

aww sweet!, cheers man, that's awesome! ;D

OpenStudy (jack1):

how'd jew find out, was it mulit choice?

OpenStudy (anonymous):

Thank you very much, Jack. I found out because when I submit the answer, it tells me if I am write or wrong. However, I only get 5 attempts. If I do not get them right those 5 times, I get nein marks.

OpenStudy (jack1):

cools, all good Hokage-sama ;D slaters dude, good luck, imma go watch the latest ep that came out this week now!

OpenStudy (anonymous):

I shall be reading the manga. Enjoy :)

OpenStudy (jack1):

dam, never started the manga, is tobi a clone of Madara?

OpenStudy (anonymous):

You want me to spoil it?

OpenStudy (jack1):

weellllllll... i guess not ;( *sigh

OpenStudy (anonymous):

I can tell you, but major spoiler xD

OpenStudy (jack1):

....nah, i guess i can wait... week to week.... ;D slaters brother!